5. To warm up some milk for a baby, a caregiver pours milk into a thin-walled glass whose diameter is 6 cm. The height of the need the solution to problem 6. I already did problem 5 --> solution to problem 5 is 5.8 minutes.
Answer #1

Solis properties of water at Arg. temp. of 20c are unit M (dysianic viscosite ) = 1002 RC Kinanutic viscoside) = lood unit SBiot mo. Bir nec Bist No. Bir hike 120x0:06 * 1000 - 12:53 3984 i = 0082 = 0+1 Messeters chart: Noves being . for Bi = 0.08NOW: Baka da 196 = 0.599 x 20:00 AT24125 R 219.25°C 5 T-20 = 21921s 33 T- 112C interface temperature will be 23825°C ca